# Instantaneous Type I blow-up and non-uniqueness of smooth solutions of the Navier–Stokes equations

### v1

@@ 1.4.2. Self-similar solutions
A different mechanism for non-uniqueness of Navier–Stokes solutions was envisioned by Jia and Šverák in [MR3179576, MR3341963]. Their program concerns forward self-similar solutions of (1.1) which evolve from (1)(-1)-homogeneous initial data. Such solutions are necessarily singular at the origin but lie in certain sharp regularity classes. More explicitly, one searches for solutions on $\mathbb{R}^{d}\times(T_{*},\infty)$ of the form
@@
If there exist two distinct profiles $U_{1},U_{2}$ with the same far-field asymptotics, then the corresponding self-similar flows yield two different smooth solutions of the Navier–Stokes equations arising from the same (1)(-1)-homogeneous initial data at $t=T_{*}$. Similarly, if there exists a profile $\overline{U}$ which is nonlinearly unstable in the time-dependent version of (1.9), then the unstable manifold is a family of solutions converging to $\overline{U}$ at the initial time $T_{*}$. This program was carried out for the forced axisymmetric Navier–Stokes equations without swirl by Albritton, Brué, and Colombo [abc], following construction of an unstable vortex for forced 2D Euler by Vishik [vishik1, vishik2]. For the unforced Navier–Stokes equations, unstable solutions of (1.9) were discovered numerically by Guillod and Šverák [MR4593274], and very recently a computer-assisted proof was announced by Hou, Wang, and Yang [hou2025nonuniqueness].
@@ 1.4.3. Non-uniqueness based on dyadic scenarios
In [CoiculescuPalasek2025], Coiculescu and Palasek introduced another mechanism for non-uniqueness, based on the construction in [palasek2024non] for the Obukhov dyadic model. The initial data lies at the critical regularity as it does for self-similar solutions, but is not precisely 1-1-homogeneous. This scenario is fundamentally different from the two above, as there is no energy cascade. Instead, it is based on an asymmetric “self-distraction” where a binary decision is made at $t=0$ at infinite wavenumber, and that information is propagated to low frequencies. The initial data is designed so that either the even frequency modes of the initial data annihilate the odd ones, or vice versa, resulting in two distinct solutions. Compare also to the recent construction [palasek2025arbitrary] of norm growth from bounded data which involves a finite inverse cascade, but with energy supplied by the large initial velocity rather than appearing anomalously from infinite frequencies.
@@ 1.5. Sharpness of the non-uniqueness theorem
The first unconditional weak-strong uniqueness theorem was obtained by Fabes, Jones, and Rivière in [MR316915], where it was proved that weak solutions are unique under the Ladyzhenskaya–Prodi–Serrin condition away from the endpoints, i.e., in the class $L^{p}_{t}L^{q}_{x}$ with $\frac{2}{p}+\frac{d}{q}\leq 1$ and $d<q<\infty$. Later, uniqueness at the endpoint $C_{t}L^{d}$ was established in [MR1813331] for $d=3$, and subsequently extended in [MR1724946, MR1680809, MR1876415] to various spatial domains using different approaches. Recently, in [2009.06596] it was shown that the method of [MR1876415] can be extended to cover the whole Ladyzhenskaya–Prodi–Serrin line including the other endpoint $L^{2}_{t}L^{\infty}_{x}$.
@@
Let $U$ be a classical solution of (1.1) on $0,T]\times\mathbb{T}^{d}$, $d\geq 2$. If $u$ is a weak solution of (1.1) with $u(0)=U(0)$ such that
$u\in L_{t}^{2}L_{x}^{\infty}([0,T]\times\mathbb{T}^{d}),$
then $u\equiv U$ for all $t\in[0,T]$.
In the case $U\equiv 0$, we have the following uniqueness conditions that follow from the energy equality for solutions satisfying some Onsager-type conditions, see [1802.05785].
@@
If $u$ is a weak solution of the (1.1) with $u(0)=0$ such that one of the following holds
@@
    $u\in L^{2}_{t}\dot{H}^{1}_{x}([0,T]\times\mathbb{T}^{d})\cap L_{t}^{\infty}B^{-1}_{\infty,\infty}([0,1]\times\mathbb{T}^{d})$,
@@
    $u\in L^{\infty}_{t}L^{2}_{x}([0,T]\times\mathbb{T}^{d})\cap L^{2}_{t}\dot{H}^{1}_{x}([0,T]\times\mathbb{T}^{d})\cap L_{t}^{2,\infty}L^{\infty}_{x}([0,T]\times\mathbb{T}^{d})$,
then $u\equiv 0$ for all $t\in[0,T]$.
While the first condition simply implies that $u\in L^{3}B^{1/3}_{3,c_{0}}$ via interpolation, the second condition requires an argument that takes advantage of the Laplacian.
@@ 1.6. Inverse energy cascade
$\Pi_{N}(I)=-\int_{I}\int_{\mathbb{T}^{d}}u\cdot\nabla u\cdot P_{N}u\,dxdt.$
@@
$\frac{1}{2}\|P_{N}u(t)\|_{L^{2}}^{2}=\frac{1}{2}\|P_{N}u(t_{0})\|_{L^{2}}^{2}+\Pi_{N}([t_{0},t])-\nu\int_{t_{0}}^{t}\|\nabla P_{N}u\|_{L^{2}}^{2}\,d\tau.$
@@
ruling out anomalous dissipation, and hence yielding the standard energy equality. In fact, the same conclusion also holds for weak solutions in the Onsager class $u\in L^{3}([0,T];B^{\frac{1}{3}}_{3,c_{0}})$, see [MR2422377], or satisfying condition of Theorem 1.11.
@@
$\lim_{N\to\infty}\Pi_{N}([T_{*},t])=\infty,\qquad\lim_{N\to\infty}\Pi_{N}([t,T])=0,$
@@
Comparison to “corner singularities” and “initial layers”. The constructed blow-up from the right in Thoerem 1.1 appears to be the first of its kind not only for fluid equations, but in the area of partial differential equations in general. The closest analog might be corner singularities. A canonical example is a heat equation with mismatched boundary and initial conditions. Consider the heat equation on the half-line $\{x>0\}$:
@@ 1.7. Main ideas of the construction
Here we give a heuristic overview of the multi-scale construction leading to Theorems 1.1 and 1.6, which, at leading order, involves infinitely-many interacting components with strong localization in time and frequency, and weak localization in space. The goal is to build a smooth solution of the Navier–Stokes equations that stays regular for $t>0$, yet whose $L^{\infty}$ norm diverges as $t\to 0^{+}$. To achieve this, we construct an infinite hierarchy of oscillatory building blocks indexed by $(j,k)\in\mathcal{J}_{d}\times\mathbb{N}$ (with $\mathcal{J}_{d}$ some finite set), each corresponding to a shell in Fourier space. The localized elements are designed so that the solution components in shells with index $k+1$ interact nonlinearly to generate those in shells with index $k$ on a particular short time scale.
@@
For $k\in\mathbb{N}$ and $j\in\mathcal{J}_{d}$ (see Section 2.1 for the definition), the frequency scales $N_{j,k}$ will be designed to rapidly increase according the ordering
@@
For a decreasing sequence of times $N_{1,k+1}^{-2}\ll t_{k}\ll N_{J_{d},k}^{-2}$, the energy transfer from the frequency levels $\{N_{j,k+1}\}_{j\in\mathcal{J}_{d}}$ to $\{N_{j,k}\}_{j\in\mathcal{J}_{d}}$ will occur on the time interval $t_{k+1},t_{k}]$; see Figure 4.
@@
Each component $v_{k}$ is generated over the short interval $t_{k-1},t_{k}]$ by self-interactions of higher modes $v_{k+1}\otimes v_{k+1}$, thereby transferring energy down the frequency ladder, producing the inverse energy cascade. To justify this ansatz, we need to show that
@@
Indeed, the separation of time scales ensures that the following integral is dominated by local interactions:
@@
for $t\geq t_{k-1}$. A precise computation showing the smallness of nonlocal interactions can be found in (4.14). Thus we have
@@
$\sum_{j}a_{j,k+1}^{2}\theta_{j}\otimes\theta_{j}=\mathcal{D}\sum_{j}N_{j,k}\psi_{j,k}+p\operatorname{Id},$
@@
$\int_{0}^{t}P_{<N_{1,k+1}}\mathbb{P}\operatorname{div}v_{k+1}\otimes v_{k+1}\,ds\approx\sum_{j}N_{j,k}\Delta\psi_{k}+\nabla p,$
@@
Figure 4. Time evolution of $\|u_{k}(t)\|_{L^{\infty}}$ and $\|u_{k+1}(t)\|_{L^{\infty}}$.
@@
To ensure that the solution is locally $L^{2}$ in space and time (which is crucial for the weak solution formulation; see Section 2.3), we additionally include a cutoff function $\varphi_{j}$ in the ansatz. (See the precise definition in Subsection 3.3 below.) As a result, the component of the solution in the frequency shell near $N_{j,k}$ occupied, to leading order, a spatial domain of volume $O(2^{-k})$.
Structure of the paper.
Section 2.1 introduces the notations to be used in the paper. Sections 3, 4 and 5 contain the technical construction and estimates. Section 6 is devoted to the proof of the main results. Auxiliary lemmas are collected in the Appendix.
Acknowledgments. M.D. is partially supported by the National Science Foundation grant DMS-2308208 and Simons Foundation. S.P. acknowledges support from the National Science Foundation under Grant No. DMS-2424441.
@@ 2.1. Notation
Let $\mathbb{T}^{d}$ denote the $d$-dimensional torus $(\mathbb{R}/2\pi\mathbb{Z})^{d}$. We define the spaces $L^{p}(\mathbb{T}^{d})$ in the standard way for $1\leq p\leq\infty$, and write the norm as $\|f\|_{L^{p}(\mathbb{T}^{d})}$ or $\|f\|_{p}$ when there is no ambiguity. We also denote by $L^{p,\infty}$ the Lorentz space (or weak $L^{p}$ space) defined in the standard way via distribution function $\lambda(s)=|\{x\in\mathbb{T}^{d}:|f(x)|>s\}|$. Here and elsewhere, $|\cdot|$ denotes the volume $\Omega$ restricted to the fundamental domain of $\mathbb{T}^{d}$. We point out the inclusion
$L^{p}\subset L^{p,\infty}$
@@
$\hat{f}(\xi)\coloneqq\fint_{\mathbb{T}^{d}}f(x)e^{-ix\cdot\xi}dx,\qquad\mathcal{F}^{-1}g(\xi)\coloneqq\sum_{\xi\in\mathbb{Z}^{d}}g(\xi)e^{ix\cdot\xi}$
for $\xi\in\mathbb{Z}^{d}$ and $x\in\mathbb{T}^{d}$, where $\fint_{\Omega}\coloneqq|\Omega|^{-1}\int_{\Omega}$ and, for $\Omega\subset\mathbb{T}^{d}$.
@@
$\mathbb{P}f\coloneqq\mathcal{F}^{-1}\left(\left(\operatorname{Id}-\frac{\xi\otimes\xi}{|\xi|^{2}}\right)\hat{f}(\xi)\right)$
@@
Let $\mathrm{Sym}^{d}(\mathbb{R})$ be the space of real symmetric $d\times d$ matrices, equipped with some norm. Let $J_{d}\coloneqq\dim\mathrm{Sym}^{d}(\mathbb{R})=d(d+1)/2$, and define $\mathcal{J}_{d}\coloneqq\{1,2,\ldots,J_{d}\}$. For two vectors $v,w\in\mathbb{R}^{d}$, we define the tensor product $(v\otimes w)_{i,j}=v_{i}w_{j}$ and symmetrized tensor product $(v\odot w)_{i,j}=\frac{1}{2}(v_{i}w_{j}+v_{j}w_{i})$. For brevity, we apply this algebraic notation to operators as well; for instance, if $f$ is a scalar function, then we denote by $\nabla\odot f$ the symmetric tensor field with components $\frac{1}{2}(\partial_{i}f_{j}+\partial_{j}f_{i})$.
@@
$\mathcal{D}f$ $\coloneqq 2\nabla\odot f-2(\operatorname{div}f)\operatorname{Id}$
$\tilde{\mathcal{D}}f$ $\coloneqq 2\nabla\odot f-(\operatorname{div}f)\operatorname{Id}$
$\mathcal{R}$ $\coloneqq\Delta^{-1}\tilde{\mathcal{D}}$
$\mathbb{Q}$ $\coloneqq 2\Delta^{-1}\nabla\odot\mathbb{P}\operatorname{div}$
@@
This version of the anti-divergence $\mathcal{R}$ is standard and improves on the original definition from [DeLellisSzekelyhidi2013] in the sense that $\Delta\mathbb{P}$ is now a classical differential operator. This is important in the sequel because it guarantees locality and $C^{1}\to L^{\infty}$ boundedness.

### v2

@@ 2.3.2. Self-similar solutions
A different mechanism for non-uniqueness of Navier–Stokes solutions was envisioned by Jia and Šverák in [27, 28]. Their program concerns forward self-similar solutions of (1.1) which evolve from (1)(-1)-homogeneous initial data. Such solutions are necessarily singular at the origin but lie in certain sharp regularity classes. More explicitly, one searches for solutions on $\mathbb{R}^{d}\times(T_{*},\infty)$ of the form
@@
If there exist two distinct profiles $U_{1},U_{2}$ with the same far-field asymptotics, then the corresponding self-similar flows yield two different smooth solutions of the Navier–Stokes equations arising from the same (1)(-1)-homogeneous initial data at $t=T_{*}$. Similarly, if there exists a profile $\overline{U}$ which is nonlinearly unstable in the time-dependent version of (2.3), then the unstable manifold is a family of solutions converging to $\overline{U}$ at the initial time $T_{*}$. This program was carried out for the forced axisymmetric Navier–Stokes equations without swirl by Albritton, Brué, and Colombo [1], following construction of an unstable vortex for forced 2D Euler by Vishik [50, 51]. For the unforced Navier–Stokes equations, unstable solutions of (2.3) were discovered numerically by Guillod and Šverák [21], and very recently a computer-assisted proof was announced by Hou, Wang, and Yang [25].
@@ 2.3.3. Non-uniqueness based on dyadic scenarios
In [12], Coiculescu and Palasek introduced another mechanism for non-uniqueness, based on the construction in [45] for the Obukhov dyadic model. The initial data lies at the critical regularity as it does for self-similar solutions, but is not precisely 1-1-homogeneous. This scenario is fundamentally different from the two above, as there is no energy cascade. Instead, it is based on an asymmetric “self-distraction” where a binary decision is made at $t=0$ at infinite wavenumber, and that information is propagated to low frequencies. The singular initial data is designed so that either the even frequency modes of the initial data annihilate the odd ones, or vice versa, resulting in two distinct solutions. Compare also to the recent construction [44] of norm growth from bounded data which involves a finite inverse cascade, but with energy supplied by the large initial velocity rather than appearing anomalously from infinite frequencies.
@@ 2.4. Sharpness of the non-uniqueness theorem
A weak solution is defined in Section 3.3 as a vector field $u\in L^{2}(\mathbb{T}^{d}\times[0,T])$ satisfying the Navier–Stokes equations in a suitable averaged sense. In [18], Fabes, Jones, and Rivière proved that weak solutions also satisfy the following integral equation, known as the mild formulation,
$u=e^{t\Delta}u_{0}-\int_{0}^{t}e^{(t-s)\Delta}\mathbb{P}\operatorname{div}(u\otimes u)(s)\,ds,$
which can be formally obtained from (1.1) using Duhamel’s principle.
The first unconditional weak–strong uniqueness theorem was also obtained in [18], where it was proved that weak solutions are unique under the Ladyzhenskaya–Prodi–Serrin condition away from the endpoints, i.e., in the class $L^{p}_{t}L^{q}_{x}$ with $\frac{2}{p}+\frac{d}{q}\leq 1$ and $d<q<\infty$. Later, uniqueness at the endpoint $C_{t}L^{d}$ was established in [19] for $d=3$, and subsequently extended in [39, 41, 37] to various spatial domains using different approaches. Recently, in [9] it was shown that the method of [37] can be extended to cover the whole Ladyzhenskaya–Prodi–Serrin line including the other endpoint $L^{2}_{t}L^{\infty}_{x}$:
@@
Let $U$ be a classical solution of (1.1) on $\mathbb{T}^{d}\times[T_{*},T]$, $d\geq 2$. If $u$ is a weak solution of (1.1) with $u(T_{*})=U(T_{*})$ such that
$u\in L_{t}^{2}L_{x}^{\infty}([T_{*},T]\times\mathbb{T}^{d}),$
then $u\equiv U$ for all $t\in[T_{*},T]$.
The non-unique solutions constructed in this paper lie at the borderline of $L^{2}_{t}L^{\infty}_{x}$, as they satisfy the Type I bound and even Orlicz-type estimates as discussed in Remark 1.3.
Another relevant uniqueness condition was proved in 2001 by Koch and Tataru (see Section 3.2 for details):
Theorem 2.2 ([32]).
The Navier–Stokes equations (1.1) have a unique small global solution in $X$ for all initial data $u_{0}$ with $\operatorname{div}u_{0}=0$ which are small in $BMO^{-1}$.
The non-unique solutions constructed here belong to the Koch–Tataru path space $X$ and are allowed to have small (even zero) initial data in $BMO^{-1}$. While the Koch–Tataru argument shows that a small ball in $X$ contains only one solution, which follows from Banach’s fixed point theorem, our non-unique solutions lie outside of this ball. More precisely, if the classical solution $U$ in Theorem 1.7 is such that $\|U(T_{*})\|_{BMO^{-1}}$ is small enough, then $U$ is a unique small Koch–Tataru solution on $T_{*},T]$, and can in fact be extended to a global classical solution on $T_{*},\infty)$. The rest of the family $(u^{(\sigma)})_{\sigma\in(0,1]}$ are of order one in the Koch–Tataru path space.
We also have the following uniqueness conditions that follow from the energy equality for weak solutions in some Onsager-type spaces.
@@
Let $U$ be a classical solution of (1.1) on $\mathbb{T}^{d}\times[T_{*},T]$, $d\geq 2$. If $u$ is a weak solution of (1.1) with $u(T_{*})=U(T_{*})$ such that one of the following holds
@@
    $u\in L^{2}_{t}\dot{H}^{1}_{x}([T_{*},T]\times\mathbb{T}^{d})\cap L_{t}^{\infty}B^{-1}_{\infty,\infty}([T_{*},T]\times\mathbb{T}^{d})$,
@@
    $u\in L^{\infty}_{t}L^{2}_{x}([T_{*},T]\times\mathbb{T}^{d})\cap L^{2}_{t}\dot{H}^{1}_{x}([T_{*},T]\times\mathbb{T}^{d})\cap L_{t}^{2,\infty}L^{\infty}_{x}([T_{*},T]\times\mathbb{T}^{d})$,
then $u\equiv U$ for all $t\in[T_{*},T]$.
Proof.
The first condition simply implies that $u\in L^{3}(T_{*},T;B^{\frac{1}{3}}_{3,c_{0}})$ via interpolation. Then the estimate on the energy flux from [7] yields the energy equality
$\tfrac{1}{2}\|u(t)\|_{L^{2}}^{2}=\tfrac{1}{2}\|u(T_{*})\|_{L^{2}}^{2}-\nu\int_{T_{*}}^{t}\|\nabla u(s)\|_{L^{2}}^{2}\,ds,\qquad T_{*}\leq t\leq T,$
which means that $u$ is a Leray–Hopf weak solution on $T_{*},T]$.
The second condition does not directly imply that $u\in L^{3}(T_{*},T;B^{\frac{1}{3}}_{3,c_{0}})$. Nevertheless, in [8], Theorem 1.1, using an argument that takes advantage of the Laplacian, it was shown that $u$ still satisfies the energy equality and hence it is a Leray–Hopf solution on $T_{*},T]$. Note that Theorem 1.1 is stated for $d=3$ only, but the proof is dimension independent.
To conclude, since $u(T_{*})$ is smooth, we can apply Leray’s weak–strong uniqueness principle, which holds for all $d\geq 2$, and says that smooth solutions are unique in the class of Leray–Hopf weak solutions. Thus $u\equiv U$ for all $t\in[T_{*},T]$.
Finally, we can show that for any weak solution exhibiting an instantaneous Type I blow-up of the $L^{\infty}$ norm, the energy must blow up as well.
Theorem 2.4.
Let $u$ be a weak solution of (1.1) on $\mathbb{T}^{d}\times[0,T]$, $d\geq 2$, such that
  1. (1)
    $u$ is a classical solution of (1.1) on $\mathbb{T}^{d}\times([0,T]\setminus\{T_{*}\})$ for some $T_{*}\in[0,T)$,
  2. (2)
    $u(T_{*})\in C^{\infty}(\mathbb{T}^{d})$,
  3. (3)
    $u$ exhibits instantaneous Type I blow-up at $T_{*}$:
    $\limsup_{t\to T_{*}+}\|u(t)\|_{L^{\infty}}=\infty,\qquad\limsup_{t\to T_{*}+}\,\sqrt{t-T_{*}}\|u(t)\|_{L^{\infty}}<\infty.$
Then the energy of $u$ instantaneously blows up at $t=T_{*}$:
$\lim_{t\to T_{*}+}\|u(t)\|_{L^{2}}=\infty.$
Proof.
Suppose to the contrary that $\|u(t)\|_{L^{2}}$ is bounded. Since $u$ is a classical solution on $(T_{*},T]$, it satisfies energy equality on every interval $t_{0},T]$ with $t_{0}\in(T_{*},T]$, and hence
$2\nu\int_{T_{*}}^{T}\|\nabla u(t)\|^{2}_{L^{2}}\,dt\leq\lim_{t\to T_{*}+}\|u(t)\|^{2}_{L^{2}}<\infty.$
Thus $u$ satisfies the second condition of Theorem 2.3 and has to coincide with a smooth solution arising from $u(T_{*})$ on some short time interval. Then
$\limsup_{t\to T_{*}+}\|u(t)\|_{L^{\infty}}<\infty,$
a contradiction. ∎
@@ 2.5. Inverse energy cascade
$\Pi_{N}(I)=-\int_{I}\int_{\mathbb{T}^{d}}(u\cdot\nabla u)\cdot P_{N}u\,dxdt.$
@@
$\tfrac{1}{2}\|P_{N}u(t)\|_{L^{2}}^{2}=\tfrac{1}{2}\|P_{N}u(t_{0})\|_{L^{2}}^{2}+\Pi_{N}([t_{0},t])-\nu\int_{t_{0}}^{t}\|\nabla P_{N}u(s)\|_{L^{2}}^{2}\,ds.$
@@
ruling out anomalous dissipation, and hence yielding the standard energy equality. In fact, the same conclusion also holds for weak solutions in the Onsager class $u\in L^{3}([0,T];B^{\frac{1}{3}}_{3,c_{0}})$, see [7], or satisfying one of the conditions of Theorem 2.3.
@@
$\lim_{N\to\infty}\Pi_{N}([T_{*},t])=+\infty,\qquad\lim_{N\to\infty}\Pi_{N}([t,T])=0,$
@@
Comparison to “corner singularities” and “initial layers”. The constructed blow-up from the right in Theorem 1.1 appears to be the first of its kind not only for fluid equations, but in the area of partial differential equations in general. The closest analog might be corner singularities. A canonical example is the heat equation with mismatched boundary and initial conditions. Consider the heat equation on the half-line $\{x>0\}$:
@@ 2.6. Main ideas of the construction
Here we give a heuristic overview of the multi-scale construction leading to Theorems 1.1 and 1.7, which, at leading order, involves infinitely many interacting components with strong localization in time and frequency, and weak localization in space. The goal is to build a smooth solution of the Navier–Stokes equations that stays regular for $t>0$, yet whose $L^{\infty}$ norm diverges as $t\to 0^{+}$. To achieve this, we construct an infinite hierarchy of oscillatory building blocks indexed by $(j,k)\in\mathcal{J}_{d}\times\mathbb{N}$ (with $\mathcal{J}_{d}$ some finite set), each corresponding to a shell in Fourier space. The localized elements are designed so that the solution components in shells with index $k+1$ interact nonlinearly to generate those in shells with index $k$ on a particular short time scale.
@@
For $k\in\mathbb{N}$ and $j\in\mathcal{J}_{d}$ (see Section 3.1 for the definition), the frequency scales $N_{j,k}$ will be designed to rapidly increase according to the ordering
@@
For a decreasing sequence of times $N_{1,k+1}^{-2}\ll t_{k}\ll N_{J_{d},k}^{-2}$, the energy transfer from the frequency levels $\{N_{j,k+1}\}_{j\in\mathcal{J}_{d}}$ to $\{N_{j,k}\}_{j\in\mathcal{J}_{d}}$ occurs on the time interval $t_{k+1},t_{k}]$; see Figure 4.
@@
Each component $v_{k}$ is generated over the short interval $t_{k+1},t_{k}]$ by self-interactions of higher modes $v_{k+1}\otimes v_{k+1}$, thereby transferring energy down the frequency ladder, producing the inverse energy cascade. To justify this ansatz, we need to show that
@@
When $d=2$, the separation of time scales ensures that the following integral is dominated by local interactions:
@@
for $t\geq t_{k-1}$. A precise computation showing the smallness of nonlocal interactions can be found in (5.14). When $d\geq 3$, these interactions can be minimized by selection of the supports of the building blocks. Thus we have
@@
$\sum_{j}a_{j,k+1}^{2}\theta_{j}\otimes\theta_{j}\approx\mathcal{D}\sum_{j}N_{j,k}\psi_{j,k}+p\operatorname{Id},$
@@
$\int_{0}^{t}P_{<N_{1,k+1}}\mathbb{P}\operatorname{div}v_{k+1}\otimes v_{k+1}\,ds\approx\sum_{j}N_{j,k}\Delta\psi_{j,k}+\nabla p,$
@@
Figure 4. Time evolution of $\|v_{k}(t)\|_{L^{\infty}}$ and $\|v_{k+1}(t)\|_{L^{\infty}}$.
@@
To ensure that the solution is locally $L^{2}$ in space and time (which is crucial for the weak solution formulation; see Section 3.3), we additionally include a cutoff function $\varphi_{j,k}$ in the ansatz. (See the precise definition in Subsection 4.3 below.) As a result, the component of the solution in the frequency shell near $N_{j,k}$ occupies, to leading order, a spatial domain of volume $O(2^{-k})$.
@@ 3.1. Notation
Let $\mathbb{T}^{d}$ denote the $d$-dimensional torus $(\mathbb{R}/2\pi\mathbb{Z})^{d}$. We define the spaces $L^{p}(\mathbb{T}^{d})$ in the standard way for $1\leq p\leq\infty$, and write the norm as $\|f\|_{L^{p}(\mathbb{T}^{d})}$ or $\|f\|_{p}$ when there is no ambiguity. We also denote by $L^{p,\infty}$ the Lorentz space (or weak $L^{p}$ space) defined in the standard way via distribution function $\lambda(s)=|\{x\in\mathbb{T}^{d}:|f(x)|>s\}|$. Here and elsewhere, $|\Omega|$ denotes the volume of $\Omega$ restricted to the fundamental domain of $\mathbb{T}^{d}$. We point out the inclusion
$L^{p}\subset L^{p,\infty},$
@@
$\hat{f}(\xi)\coloneqq\fint_{\mathbb{T}^{d}}f(x)e^{-ix\cdot\xi}dx,\qquad\mathcal{F}^{-1}g(x)\coloneqq\sum_{\xi\in\mathbb{Z}^{d}}g(\xi)e^{ix\cdot\xi},$
for $\xi\in\mathbb{Z}^{d}$ and $x\in\mathbb{T}^{d}$, where $\fint_{\Omega}\coloneqq|\Omega|^{-1}\int_{\Omega}$ for $\Omega\subset\mathbb{T}^{d}$.
@@
$\mathbb{P}f\coloneqq\mathcal{F}^{-1}\left(\left(\operatorname{Id}-\frac{\xi\otimes\xi}{|\xi|^{2}}\right)\hat{f}(\xi)\right),$
@@
Let $\mathrm{Sym}^{d}(\mathbb{R})$ be the space of real symmetric $d\times d$ matrices, equipped with some norm. Let $J_{d}\coloneqq\dim\mathrm{Sym}^{d}(\mathbb{R})=d(d+1)/2$, and define $\mathcal{J}_{d}\coloneqq\{1,2,\ldots,J_{d}\}$. For two vectors $v,w\in\mathbb{R}^{d}$, we define the tensor product $(v\otimes w)_{i,j}=v_{i}w_{j}$ and symmetrized tensor product $(v\odot w)_{i,j}=\frac{1}{2}(v_{i}w_{j}+v_{j}w_{i})$. For brevity, we apply this algebraic notation to operators as well; for instance, if $f$ is a vector function, then we denote by $\nabla\odot f$ the symmetric tensor field with components $\frac{1}{2}(\partial_{i}f_{j}+\partial_{j}f_{i})$.
@@
$\mathcal{D}f$ $\coloneqq 2\nabla\odot f-2(\operatorname{div}f)\operatorname{Id},$
$\tilde{\mathcal{D}}f$ $\coloneqq 2\nabla\odot f-(\operatorname{div}f)\operatorname{Id},$
$\mathcal{R}$ $\coloneqq\Delta^{-1}\tilde{\mathcal{D}}$,
$\mathbb{Q}$ $\coloneqq 2\Delta^{-1}\nabla\odot\mathbb{P}\operatorname{div},$
@@
This version of the anti-divergence $\mathcal{R}$ is standard and improves on the original definition from [35] in the sense that $\Delta\mathbb{P}$ is now a classical differential operator. This is important in the sequel because it guarantees locality and $C^{1}\to L^{\infty}$ boundedness.